OpenStudy (anonymous):

One leg of a right triangle has length 7 and all sides are whole numbers. Find the lengths of the other two sides.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} x^2=y^2+7^2 \hspace{.33em}\\~\\ x^2-y^2=49 \hspace{.33em}\\~\\ (x-y)(x+y)&=49\times 1 \hspace{.33em}\\~\\ (x-y)(x+y)&=1\times 49 \hspace{.33em}\\~\\ (x-y)(x+y)&=7\times 7 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

as they are whole numbers only three types of system of equations are possible \(\large \color{black }{\begin{align} (x-y)=49 \hspace{.33em}\\~\\ (x+y)=1\hspace{.33em}\\~\\ \end{align}}\) or \(\large \color{black }{\begin{align} (x-y)=1 \hspace{.33em}\\~\\ (x+y)=49\hspace{.33em}\\~\\ \end{align}}\) or \(\large \color{black }{\begin{align} (x-y)=7 \hspace{.33em}\\~\\ (x+y)=7\hspace{.33em}\\~\\ \end{align}}\)